Metronidazole is used in the treatment of anaerobic protozoal infections. Although often administered orally, it can also be delivered by means of an IV infusion. The loading dose is 15 mg per kilogram of body mass and is infused over one hour. The maintenance dose is 7.5 mg per kilogram of body mass, is also infused over an hour, and is administered six hours after the loading dose. If the concentration of the metronidazole solution is 5 mg/ml, how many milliliters are required for a loading dose and three maintenance doses for a patient who weighs 220 pounds?
Answer #1

Loading dose 15mg /kg

Maintenance dose 7.5mg/kg

Concentration of the medicine 5mg/ml

Weight of the patient 220 pounds = 99.7903kg (1pound = 0.453592kg)

So, loading dose for this patient 15* 99.79 = 1496.85mg

The solution should include 1496.85/ 5 ml = 299.37ml

Maintenance dose = 7.5mg/kg = 7.5* 99.79 = 748.425ml

The solution needed for the maintenance dose= 748.425/5 = 149.685ml